OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using substitution: 2x − 4y = 18 3x − y + 10 = 0 Multiply the second equation by -4 Write the second equation in terms of y Multiply the first equation by 2 Replace the right side of the first equation with the second equation

OpenStudy (anonymous):

Write the second equation in terms of y. If you do that, it's then y = something, and you can SUBSTITUTE that in the first equation (that's why it's called substitution)

OpenStudy (anonymous):

Solve using substitution: -y − 5x = -7 2y − 2x = 14 Infinitely many solutions No solution (7/5, 14) (0, 7)

OpenStudy (anonymous):

Okay, multiply the first equation by 2 so now you have -2y -10x = -14 2y -2x = 14 (add together, y's and 14's cancel out) -12x = 0 x = 0 sub back in to get y 2y - 0 = 14 2y = 14 y =7 therefore, the solution is (0, 7)

OpenStudy (anonymous):

Solve using substitution: -2x − y + 1 = 0 2x = -13 − 3y (1, -9) (3, -5) (-2, 5) (4, -7) WHAT ABOUT THIS ONE

OpenStudy (anonymous):

okay, for this one, just move -y over to the other side in the first equation so now you have y = 1 - 2x sub that into the second equation 2x = -13 -3y 2x = -13 - 3(1-2x) 2x = -13 -3 +6x 2x = -16 + 6x -4x = -16 x = 4 and then find y... y = 1 - 2x y = 1 - 2(4) y = 1 - 8 y = -7 so your final answer is (4, -7)